certain reaction has an activation energy of 61.0 kJ/mol and a frequency factor of A1 = 8.90×1012 M−1s−1 . What is the rate constant, k, of this reaction at 23.0 ∘C ?
Answer #1

We use the arrhenius equation to calculate the rate constant,

K=Ae^{\frac{-Ea}{RT}}

K=8.90*10^{12}M^{-1}s^{-1}e^{\frac{-61000J/mol}{8.314J/molK*296K}}=153.12M^{-1}s^{-1}